if an organism heterozygous at two loci is crossed with an individual that is heterozygous for one locus and homozygous dominant for the other, what genotypes can be expected and at what frequency.
Answer #1

Ans. Suppose in a plant red flower(R) is dominant over white flower (r) and at other locus yellow seeds are dominant over green seeds.

So, according to the question :-

Heterozygous at two loci crossed with heterozygous at one locus and homozygous dominant for the other

RrYy X RrYY

Gametes RY, Ry, rY, ry x RY, rY

F1 progeny = RRYY, RRYy, RrYY, RrYy

RrYY, RrYy, rrYY, rrYy

RRYY = 1/8

RRYy =1/8

RrYY = 2/8

RrYy =2/8

rrYY =1/8

rrYy= 1/8